If x^2+y^2=58 and x+y=10 .find the value of x and y

Grade:11

1 Answers

Sai
5 years ago
x^2+y^2=58______eq(1)x+y=10 => y=10-x___eq(2)Sub (2) in (1),x^2+(10-x)^2=58x^2+100+x^2-20x=582x^2-20x+42=0x^2-10x+21=0(x-7)(x-3)=0x=3 (or) 7If x=3,y=10-x=10-3=7 [from (2)]If x=7,y=10-x=10-7=3 from (2)]So x&y are 3&7 (or) 7&3.